# Estimating Correctness Without Oracles in LLM-Based Code Generation

### v1

@@ Abstract
Generating code from natural language specifications is one of the most successful applications of Large Language Models (LLMs). Yet, they hallucinate: LLMs produce outputs that may be grammatically correct but are factually incorrect. Without an existing, correct implementation (i.e., an oracle), can we quantify how likely the generated program is correct?
In this paper, we propose a measure of incorrectness, called incoherence, that can be estimated efficiently in the absence of an oracle and provides a lower bound on the error, i.e., the probability that the LLM-generated program for that specification is incorrect. Our experiments demonstrate an extraordinary effectiveness. For the average code generation task, our incoherence-based methodology can automatically identify about two-thirds of incorrect programs without reports of false positives. In fact, an oracle-based evaluation of LLMs can be reliably replaced by an incoherence-based evaluation. In particular, we find a very strong agreement between the ranking of LLMs by the number of programs deemed correct via an oracle and the ranking of LLMs by the number of programs deemed correct via incoherence.
@@ 1 Introduction
LLMs have demonstrated remarkable performance on code generation tasks. Yet, confabulation remains a key concern. Models often produce syntactically correct but functionally incorrect code, raising the critical question of when such outputs can be trusted. For instance, Fan et al. (2023) found that the vast majority of auto-generated programs for easy to medium LeetCode programming tasks are incorrect and explain that 57% of those do not even properly implement the task (“algorithmic misalignment”) while 19% can only be fixed by changing multiple different code locations (multi-hunk). Pearce et al. (2025) analyzed code generated in scenarios relevant to high-risk cybersecurity weaknesses and found that 40% of the 1.7k LLM-generated programs actually contain security vulnerabilities.
@@
Our work continues a recent stream of works addressing the confabulation problem using the disagreement between independently sampled responses to detect untruthful or erroneous outputs (Manakul, Liusie, and Gales 2023; Friel and Sanyal 2023; Li et al. 2024; Farquhar et al. 2024). A high disagreement indicates a high factual inconsistency.
@@
[…]
@@
In this work, we formalize this intuition. We argue that incoherence—measured as behavioral divergence across samples—is a principled and theoretically justified proxy for model error. Concretely, given an LLM and a programming task $d$, we call the probability that any two programs generated to implement $d$ are functionally different as the LLM’s incoherence on $d$. If we are also given a ground truth implementation $f^{\ast}_{d}$ for $d$, we call the probability that $f^{\ast}_{d}$ and a program generated to implement $d$ are functionally different as the LLM’s error on $d$. We note that the widely-used pass@1 score (Chen et al. 2021; Liu et al. 2023) on a benchmark set can be written in terms of the empirical error (i.e., the error’s maximum likelihood estimate) on each task.
@@
[…]
In summary, our paper makes the following contributions:
@@
    We propose the first formal (rather than heuristic) unsupervised measure of LLM correctness, called incoherence, that can be used to estimate an LLM’s error on a coding task in the absence of an oracle.
@@
    We develop a formal probabilistic framework showing that incoherence provides a provable lower bound on the model’s error, that incoherence can be estimated efficiently with PAC-style guarantees, and how the widely-used pass@1 is formally linked to the model’s error.
@@
    We perform a large-scale empirical study involving 16 state-of-the-art LLMs on two standard code generation benchmarks, which shows that incoherence alone can detect two-thirds of incorrect generations without false positives and yields rankings of LLMs that strongly agree with an oracle-based evaluation, thus providing a reliable and scalable alternative to pass@1.
@@
    We publish prototype, results, and analysis scripts at https://github.com/mpi-softsec/difftrust.
@@ 2 Background
Code generation is by far the most prominent use case of LLMs in software engineering, according to a recent survey (Hou et al. 2024). Only four years ago, Copilot [^1] https://copilot.microsoft.com/ was one of the first widely available LLM-based code generators. Today, it has already gathered more than 40 million installations. While many other code-specific LLMs have since been developed, general-purpose LLMs are turning out to be effective coders in their own right (Hou et al. 2024; Leaderboards 2025). Indeed, in June 2025, about a quarter of the 7.5 trillion tokens passing through OpenRouter [^2] OpenRouter offers a unified API to all popular LLMs:
https://openrouter.ai/rankings/programming?view=month
as prompts to various LLMs were related to programming. Jiang et al. (2024) provide an excellent survey of recent work in LLM-based code generation and point out that trustworthiness is critical for a wide and continued adoption. LLMs are prone to generating incorrect code. So, how can we ensure that the generated code correctly solves the programming task?
Recent efforts to evaluate factual reliability (a.k.a. confabulation) in text generation have focused on measuring internal consistency across model outputs. Approaches such as SelfCheckGPT (Manakul, Liusie, and Gales 2023) and ChainPoll (Friel and Sanyal 2023) sample multiple responses from an LLM and identify inconsistencies as indicators of potential errors. (Farquhar et al. 2024) propose the notion of semantic entropy to identify confabulations. Similar LLM-generated free-form outputs are clustered (where similarity is established by an LLM), and entropy is computed from the distribution over these clusters. This idea has also been adapted to code generation. For instance, HonestCoder (Li et al. 2024) uses multiple modalities (e.g., syntax, data flow) to establish similarity among the sampled programs and to construct clusters. (Sharma and David 2025) propose to use functional equivalence between programs to construct clusters, where difference-revealing inputs are sought using symbolic execution.
Despite their strong empirical performance, these methods remain fundamentally heuristic. Their confidence estimates rely on various forms of answer comparison—such as representation-level similarity, entailment approximations, LLM prompting, or chain-of-thought reasoning—none of which can be externally validated against a definitive ground truth. In structured domains like code generation, these techniques are especially limited. Methods like chain-of-thought reasoning, as used in state-of-the-art systems such as ChainPoll (Friel and Sanyal 2023), struggle to capture the full semantic complexity of programs. This limitation is precisely what necessitates the use of correctness proxies: LLMs often fail to account for the intricate syntax, control flow, and type-level semantics that determine whether code is correct.
The problem of automatically determining whether the output of a program is correct for a given input is called oracle problem in software testing (Barr et al. 2015). Some behaviour, such as crashes or memory corruption, is unambiguously bad and can be detected by sanitizers (Serebryany et al. 2012). All other solutions are domain-specific.
@@ 3 Defining Error and Incoherence
We consider the task of automatically generating a function implementation from a natural language specification. Formally, given a textual description $d$ of a programming task, a code generation system $\mathsf{Coder}$—treated as a black-box stochastic process—samples a program $\pi\sim\mathsf{Coder}(d)$ intended to satisfy the task description $d$. Our objective is to assess the correctness of these implementations $\mathsf{Coder}(d)$ without supervision, reference solutions, or access to model internals.
@@ 3.1 Notation
We consider a probability space $(\Omega,\mathcal{P}(\Omega),\mathbb{P})$ where $\Omega$ is the sample space, $\mathcal{P}(\Omega)$ is the event space, and $\mathbb{P}$ is the probability function. The set of random variables taking values in a set $V$ consists of functions from $\Omega$ to $V$, which we denote as $V^{\Omega}$. For any mathematical expression $expr[X_{1},\dots,X_{n}]$ involving random variables $X_{1},\dots,X_{n}$, we denote by $\{expr[X_{1},\dots,X_{n}]\}$ the event $\{\omega\in\Omega\text{ }|\text{ }expr[X_{1}(\omega),\dots,X_{n}(\omega)]\}$. To simplify notation, we write $\mathbb{P}(expr)$ instead of ($\mathbb{P}(\{expr\})$. We refer to $\mathbb{E}$ as the expectation and to $\mathbb{I}_{expr}$ as the indicator function of $expr$.
@@ 3.2 Error of a Code Generation System
We model a code generation system $\mathsf{Coder}$ as a function that maps each task $d\in\mathsf{Descr}$ to a corresponding probability distribution over $\mathsf{Prog}$. Formally, for all $d\in\mathsf{Descr}$ :
@@
For every description $d\in\mathsf{Descr}$, we assume there exist an input set $\mathsf{Input}_{d}$, an output set $\mathsf{Output}_{d}$ and a correct ground truth implementation $\pi^{\ast}_{d}\in\mathsf{Prog}$ with its functional interpretation $f^{\ast}_{d}:=\llbracket\pi^{\ast}_{d}\rrbracket$ such that $f^{\ast}_{d}:\mathsf{Input}_{d}\to\mathsf{Output}_{d}$.
The pass@1 score (Chen et al. 2021) is a standard metric to evaluate the performance of $\mathsf{Coder}$. For a finite set of tasks $S\subset\mathsf{Descr}$, pass@1 is defined as the expected fraction of sampled programs that are functionally equivalent to the ground truth implementation:
[…]
@@
Moreover, we introduce a probabilistic interpretation of correctness with respect to (w.r.t.) a distribution of inputs. Rather than asking whether the generated function is correct for all inputs, which is undecidable due to Rice’s theorem, we ask whether it is correct for a typical input, drawn from a distribution that represents expected usage. We use this probabilistic interpretation of correctness to introduce a pointwise notion of the error such that a non-zero pointwise error implies a non-zero functional error (as defined in Equation (4)).
We model an input generation system $\mathsf{Gen}$ as a function that maps each task $d\in\mathsf{Descr}$ to an intuitively realistic probability distribution over the corresponding input set $\mathsf{Input}_{d}$. Formally, for all $d\in\mathsf{Descr}$ :
@@
While the functional error can be computed only by verification of functional equivalence (an undecidable problem), the pointwise error can be estimated efficiently (c.f. Appendix B.1)—in the presence of the oracle $f^{\ast}_{d}$.
@@
Our pointwise error $\operatorname{\mathcal{E}}_{\mathsf{Gen}}(d)$ models the practical reality that a program might be correct on almost all inputs that are empirically observed when the program is tested, deployed, or used in practice. By evaluating the probability of failure on a representative input distribution, the pointwise error provides a meaningful and practical estimate of the model’s reliability in practical scenarios. The pointwise error also formalizes the experimental setup originally proposed and now widely used to estimate pass@1 (i.e., the complement of the mean functional error on a fixed set of programming tasks) using a fixed set of random test cases (Chen et al. 2021; Liu et al. 2023).
@@ 4 Incoherence of a Code Generation System
Our core challenge is to estimate the pointwise error $\operatorname{\mathcal{E}}_{\mathsf{Gen}}(d)$ in the absence of the oracle $f^{\ast}_{d}$, i.e., without supervision. We aim to achieve this using only observations from sampled implementations, without relying on any internal details of $\mathsf{Coder}$. To this end, we specialize the disagreement-based hallucination detection approach (Manakul, Liusie, and Gales 2023) to the domain of code generation. The precise definition of the (probabilistic) correctness of a program w.r.t. an oracle (i.e., a ground truth implementation) provides us with the unique opportunity to formalize the approach and to introduce actual probabilistic guarantees.
@@
$\operatorname{\mathcal{I}}_{\mathsf{Gen}}(d):=\mathbb{P}\left(\llbracket\Pi^{d}_{1}\rrbracket(X)\neq\llbracket\Pi^{d}_{2}\rrbracket(X)\right).$ (8)
where $\Pi^{d}_{1},\Pi^{d}_{2}\overset{iid}{\sim}\mathsf{Coder}(d)$ are two independently sampled programs and $X\sim\mathsf{Gen}(d)$ is an input sampled from $\mathsf{Gen}$ for task $d$. This quantity captures the model’s internal uncertainty as revealed through behavioral divergence. Crucially, $\operatorname{\mathcal{I}}_{\mathsf{Gen}}(d)$ is fully observable and efficient to estimate without an oracle (see Appendix B.2).
In the following, we show that this notion of pointwise incoherence provides a rigorous lower bound on the pointwise error and that it can be efficiently estimated. In Appendix C, we develop the notion of functional incoherence and establish functional incoherence as a lower bound on the functional error in parallel.
4.1 Incoherence as Lower Bound on Error
@@
See Appendix A. ∎
This result establishes $\operatorname{\mathcal{I}}_{\mathsf{Gen}}(d)$ as a sound and theoretically grounded proxy for estimating model error on $d$. Unlike heuristic confidence scores or divergence metrics based on representation-level similarity, $\operatorname{\mathcal{I}}_{\mathsf{Gen}}(d)$ directly, precisely, and formally captures observable functional disagreement.
@@ 4.2 Incoherence is Efficiently Estimated
There exists a randomized algorithm that, given parameters $\delta>0$, $\epsilon>0$, code generator $\mathsf{Coder}$, input generator $\mathsf{Gen}$, and task $d\in D$, computes ¯$\bar{\operatorname{\mathcal{I}}}_{\mathsf{Gen}}(d)$ such that (|¯$\mathbb{P}(|\bar{\operatorname{\mathcal{I}}}_{\mathsf{Gen}}(d)-\operatorname{\mathcal{I}}_{\mathsf{Gen}}(d)|\leq\epsilon)\geq 1-\delta$ using at most $\left\lceil\frac{\log(2/\delta)}{2\epsilon^{2}}\right\rceil$ samples.
A similar theorem for the pointwise error, the randomized algorithms using Monte Carlo estimation, and the proofs for both theorems using a trivial application of Höffdings inequality are postponed to Appendix B.1B.2.
@@
A randomized algorithm based on Monte Carlo estimation and the proof is provided in Appendix B.3.

### v2

# Incoherence as Oracle-less Measure of Error in LLM-Based Code Generation

@@ Abstract
Generating code from a natural language programming task is one of the most successful applications of Large Language Models (LLMs). Yet, the generated program may be buggy. Without an oracle, such as an existing, correct implementation or a formal specification, can we somehow estimate how likely the generated program is correct?
In this paper, we propose a measure of incorrectness, called incoherence, that can be estimated efficiently in the absence of an oracle and allows us to establish a lower bound on the error, i.e., the probability that the LLM-generated program for that specification is incorrect. In our experiments, our incoherence-based methodology can automatically identify about two-thirds of incorrect programs without reports of false positives for the average task. In fact, an oracle-based evaluation of LLMs can be reliably replaced by an incoherence-based evaluation. In particular, we find a very strong agreement between the ranking of LLMs by the number of programs deemed correct via an oracle (pass@1) and the ranking of LLMs by the number of programs deemed correct via incoherence.
Data & analysishttps://github.com/mpi-softsec/difftrust
Extended versionhttps://arxiv.org/abs/2507.00057
@@ 1 Introduction
LLMs have demonstrated remarkable performance on code generation tasks. Yet, confabulation remains a key concern. Models often produce syntactically correct but functionally incorrect code, raising the critical question of when such outputs can be trusted. For instance, bugsLLM2 found that the vast majority of auto-generated programs for easy to medium LeetCode programming tasks are incorrect and explain that 57% of those do not even properly implement the task (“algorithmic misalignment”) while another 19% can only be fixed by changing multiple different code locations (multi-hunk). bugsLLM analyzed code generated in scenarios relevant to high-risk cybersecurity weaknesses and found that 40% of the 1.7k LLM-generated programs actually contain security vulnerabilities.
@@
Our work continues a recent stream of works addressing the confabulation problem using the disagreement between independently sampled responses to detect untruthful or erroneous outputs (manakul2023selfcheckgpt; friel2023chainpoll; li2024honestcoder; farquhar2024semantic). A high disagreement indicates a high factual inconsistency.
@@
Figure 1: Rankings of 16 LLMs on the two most popular code generation benchmarks, MBPP and HumanEval. “Rank 1” indicates the highest probability of producing correct programs. X-axis: Ranking in terms of $1-\texttt{pass@1}]$ (i.e., the proportion of tasks with non-zero empirical error). Y-axis: Ranking in terms of the proportion of tasks with non-zero empirical incoherence. Note that incoherence can be estimated in the absence of a ground truth implementation.
@@
In this work, we formalize this intuition. We argue that incoherence, i.e., the behavioral divergence across samples, is a principled and theoretically justified proxy for model error. Concretely, given an LLM and a programming task $d$, we call the probability that any two programs generated to implement $d$ are functionally different as the LLM’s incoherence on $d$. If we are also given a ground truth implementation $f^{\ast}_{d}$ for $d$, we call the probability that $f^{\ast}_{d}$ and a program generated to implement $d$ are functionally different as the LLM’s error on $d$. We note that the widely-used pass@1 score (chen2021evaluating; EvalPlus) on a benchmark set can be written in terms of the empirical error (i.e., the error’s maximum likelihood estimate) on each task.
@@
In experiments with 16 state-of-the-art LLMs and two popular code generation benchmarks, our incoherence measure, which requires no ground truth implementation, works incredibly well as a substitute for pass@1. Figure 1 shows rankings of those LLMs, both in terms of the proportion of tasks with non-zero empirical error (i.e., $1-\texttt{pass@1}]$) and the proportion of tasks with non-zero empirical incoherence. These rankings very strongly agree despite the absence of oracles for our incoherence measure ($\rho\geq 0.92$). We also find that a non-zero incoherence effectively detects about two-thirds of the non-zero errors in the absence of a ground truth implementation (69% and 66% detection rate on MBPP and HumanEval, resp.). No false positives. In cases where the incoherence is zero, the mean error is substantially lower than the average. If we increase the number of generated programs for a programming task 5-fold (from 10 to 50), the detection rate further increases by eight (8) percentage points—of course, at the cost of a 5-fold increase in monetary expenses for additional queries to the LLM.
Contributions:
@@
    We propose incoherence, a formal, unsupervised proxy for correctness that lower-bounds LLM error.
@@
    We develop a probabilistic framework linking incoherence to pass@1, with PAC-style efficiency guarantees.
@@
    Our study shows, incoherence detects errors reliably and yields rankings that agree with oracle-based evaluations.
@@
    We release all code, results, and analysis scripts.
@@ 2 Background
Code generation is a primary application of LLMs in software engineering (llm4SEsurvey). Tools like Copilot have seen widespread adoption, with over 40 million installations. Today, even general-purpose LLMs are competitive coders (leaderboards). In June 2025, nearly 25% of the 7.5 trillion tokens on (openrouter) were coding-related.
llm4codeSurvey highlight that trustworthiness is essential for LLM adoption in programming. Yet, LLMs are prone to generating incorrect code. How can correctness be validated without ground truth? Existing work addresses factuality in text by sampling multiple responses and measuring internal consistency—e.g., SelfCheckGPT (manakul2023selfcheckgpt), ChainPoll (friel2023chainpoll), and semantic entropy approaches (farquhar2024semantic). These methods cluster outputs and estimate confidence based on entropy or token overlap. For code generation, variants include HonestCoder (li2024honestcoder), which uses syntax and data-flow modalities, and functional equivalence methods using symbolic execution (sharma2025assessingcorrectnessllmbasedcode). However, these approaches are fundamentally heuristic and lack formal guarantees. Their reliance on patterns or output similarity do not capture semantic correctness.
corina are interested in generating the most likely correct program for a task, and propose to select from a pool of candidates generated by multiple LLMs that program which best aligns with consensus, where consensus is defined both syntactically and semantically.
Our work addresses the oracle problem in software testing (oracle). While runtime crashes are detectable via sanitizers (asan), functional correctness is domain-specific and often unobservable. We address this problem by viewing the generated program as a random variable and using incoherence as a proxy for correctness.
@@ 3 Defining Error and Incoherence
We consider the task of automatically generating a program from a natural language specification. Formally, given a textual description $d$ of a programming task, a code generation system $\mathsf{Coder}$—treated as a black-box stochastic process—samples a program $\pi\sim\mathsf{Coder}(d)$ intended to satisfy the task description $d$. Our objective is to assess the correctness of these programs $\mathsf{Coder}(d)$ without supervision, reference solutions, or access to model internals.
@@ 3.1 Notation
We write $\mathbb{P}(\cdot)$ for probability and $\mathbb{E}[\cdot]$ for expectation, leaving the underlying probability space implicit. For any expression expr involving random variables, $\mathbb{P}(\textit{expr})$ denotes the probability of the corresponding event. We use $\mathbb{I}(\textit{expr})$ for the indicator function, which is $1$ when expr holds and 0 otherwise.
@@ 3.2 Error of a Code Generation System
We model a code generation system $\mathsf{Coder}$ as a function that maps each task $d\in\mathsf{Descr}$ to a corresponding (unknown) distribution over $\mathsf{Prog}$. Formally, for all $d\in\mathsf{Descr}$ :
@@
For every task description $d\in\mathsf{Descr}$, we assume there exist an input set $\mathsf{Input}_{d}$, an output set $\mathsf{Output}_{d}$ and a correct (deterministic) ground truth implementation $\pi^{\ast}_{d}\in\mathsf{Prog}$ with its functional interpretation $f^{\ast}_{d}:=\llbracket\pi^{\ast}_{d}\rrbracket$ such that $f^{\ast}_{d}:\mathsf{Input}_{d}\to\mathsf{Output}_{d}$.
The pass@1 score (chen2021evaluating) is a standard metric to evaluate the performance of $\mathsf{Coder}$. For a finite set of tasks $S\subset\mathsf{Descr}$, pass@1 is defined as the expected fraction of sampled programs that are functionally equivalent to the ground truth implementation:
$\text{{pass@1}}(S):=\mathbb{E}\left[\frac{1}{|S|}\sum_{d\in S}\mathbb{I}\bigl(\llbracket\Pi^{d}\rrbracket=f^{\ast}_{d}\bigr)\right].$ (3)
@@
Moreover, we introduce a probabilistic interpretation of correctness with respect to (w.r.t.) a distribution of inputs. Rather than asking whether the generated function is correct for all inputs, which is undecidable due to Rice’s theorem, we ask whether it is correct for a typical input, drawn from a distribution that represents expected usage. We use this probabilistic interpretation of correctness to introduce a pointwise notion of the error such that a non-zero pointwise error implies a non-zero functional error (cf. Eq. (4)).
We model an input generation system $\mathsf{Gen}$ as a function that maps each task $d\in\mathsf{Descr}$ to an (unknown) probability distribution over the corresponding input set $\mathsf{Input}_{d}$. This distribution might represent how the program is executed under a typical workload. Formally, for all $d\in\mathsf{Descr}$ :
@@
While the functional error can be computed only by verification of functional equivalence (an undecidable problem), the pointwise error can be estimated efficiently (c.f. Appendix C.1)—in the presence of the oracle $f^{\ast}_{d}$.
@@
Our pointwise error $\operatorname{\mathcal{E}}_{\mathsf{Gen}}(d)$ models the practical reality that a program might be correct on almost all inputs that are empirically observed when the program is tested, deployed, or used in practice. By evaluating the probability of failure on a representative input distribution, the pointwise error provides a meaningful and practical estimate of the model’s reliability in practical scenarios. The pointwise error also formalizes the experimental setup originally proposed and now widely used to estimate pass@1 (i.e., the complement of the mean functional error on a fixed set of programming tasks) using a fixed set of random test cases (chen2021evaluating; EvalPlus).
@@ 4 Incoherence of a Code Generation System
Our core challenge is to estimate the pointwise error $\operatorname{\mathcal{E}}_{\mathsf{Gen}}(d)$ in the absence of the oracle $f^{\ast}_{d}$, i.e., without supervision. We aim to achieve this using only observations from sampled implementations, without relying on any internal details of $\mathsf{Coder}$. To this end, we specialize the disagreement-based hallucination detection approach (manakul2023selfcheckgpt) to the domain of code generation. The precise definition of the (probabilistic) correctness of a program w.r.t. an oracle (i.e., a ground truth implementation) provides us with the unique opportunity to formalize the approach and to introduce actual probabilistic guarantees.
@@
$\operatorname{\mathcal{I}}_{\mathsf{Gen}}(d):=\mathbb{P}\left(\llbracket\Pi^{d}_{1}\rrbracket(X)\neq\llbracket\Pi^{d}_{2}\rrbracket(X)\right)\\$ (8)
where $\Pi^{d}_{1},\Pi^{d}_{2}\overset{iid}{\sim}\mathsf{Coder}(d)$ are two independently sampled programs and $X\sim\mathsf{Gen}(d)$ is an input sampled from $\mathsf{Gen}$ for task $d$. This quantity captures the model’s internal uncertainty as revealed through behavioral divergence. Crucially, $\operatorname{\mathcal{I}}_{\mathsf{Gen}}(d)$ is fully observable and efficient to estimate without an oracle (see Appendix C.2).
In the following, we show that this notion of pointwise incoherence provides a rigorous lower bound on the pointwise error and that it can be efficiently estimated. In Appendix D, we develop the notion of functional incoherence and establish a lower bound on the functional error in terms of the functional incoherence in parallel.
4.1 Lower Bound on Error in Terms of Incoherence
@@
See Appendix B. ∎
This result establishes $\operatorname{\mathcal{I}}_{\mathsf{Gen}}(d)$ as a sound and theoretically grounded proxy for estimating model error on $d$. Unlike heuristic confidence or divergence metrics based on representation-level similarity, $\operatorname{\mathcal{I}}_{\mathsf{Gen}}(d)$ directly, precisely, and formally captures observable functional disagreement.
@@ 4.2 Incoherence is Efficiently Estimated
There exists a randomized algorithm that, given parameters $\delta>0$, $\epsilon>0$, code generator $\mathsf{Coder}$, input generator $\mathsf{Gen}$, and task $d\in D$, computes $\bar{\operatorname{\mathcal{I}}}_{\mathsf{Gen}}(d)$ such that $\mathbb{P}(|\bar{\operatorname{\mathcal{I}}}_{\mathsf{Gen}}(d)-\operatorname{\mathcal{I}}_{\mathsf{Gen}}(d)|\leq\epsilon)\geq 1-\delta$ using at most $\left\lceil\frac{\log(2/\delta)}{2\epsilon^{2}}\right\rceil$ samples.
A similar theorem for the pointwise error, the randomized algorithms using Monte Carlo estimation, and the proofs for both theorems using a trivial application of Hoeffdings inequality are postponed to Appendix C.1C.2.
@@
A randomized algorithm based on Monte Carlo estimation and the proof is provided in Appendix C.3.